The electrical breakdown in air occurs at electric field strength of 3.5MV m−1. Calculate the corresponding intensity of a plane wave
1
Expert's answer
2019-03-13T12:47:33-0400

The intensity of a plane wave


"I=\frac{c\varepsilon_0E^2}{2}"

"=\frac{3\times 10^8\:\rm{m/s}\times 8.85\times 10^{-12}\:\rm{F/m}\times (3.5\times 10^6\:\rm{V/m})^2}{2}"

"=1.63\times 10^{10}\:\rm{W/m^2}"